OpenStudy (anonymous):

The function f(x)=9x+2^x−1 has one local minimum and one local maximum. This function has a local minimum at x equals with value and a local maximum at x equals with value

OpenStudy (anonymous):

DO U HAVE ANSWER CHOICES

OpenStudy (anonymous):

no they didnt give me answer choices

OpenStudy (anonymous):

this f(x)=9x+2x−1 is a line u forget one exponent i think

OpenStudy (anonymous):

let me rewrite it is 9x+2x^-1

OpenStudy (anonymous):

again find the derivative and the n the critical points

OpenStudy (stamp):

\[f(x)=9x+2x^{-1},\ f(x)=9x+2/x\]\[f'(x)=9-2/x^2\]Find x when f'(x) = 0\[f'(x)=9-2/x^2\]\[0=9-2/x^2\]\[x^2=9/2\]\[x=±\sqrt2/3\]

OpenStudy (stamp):

Test it for max and min at the positive x and negative x values